# 61A-Bot: AI homework assistance in CS1 is fast and cheap—but is it helpful?

### v1

@@ 4.2. Student Reception
We sought student feedback in two ways. First, we queried students for their assessment of each individual hint, which we received for approximately 27% () of queries.666These counts reflect queries from only those students who gave consent for their data to be used for research. Of these, 70% () were rated as “helpful,” with 45% of these, , reporting that the problem was now resolved. A further 10% () were rated as “not helpful, but made sense,” while the final 20% () were rated as insensible, misleading, or wrong.
@@
In Spring 2024, we conducted a non-anonymous survey at the end of the semester, to which 89% () students responded. On a scale from 1 to 5, we asked students to rate their bot usage, bot helpfulness, bot reliability, and overall satisfaction with the bot. Finally, we asked them whether or not they recommend that the bot be available to students in future semesters. The results from these surveys can be found in Figure 2.
@@
In the Spring 2024 survey, we also asked two optional questions: What did students like the most and least about learning with the bot? We include a representative response for each here: What students liked the most:
@@
“What I loved about the bot is that it allowed me to get feedback when I didn’t have access to a tutor. Accordingly, instead of banging my head against the wall for hours, I was able to get feedback about what I was doing wrong and correct the mistakes. For me personally, I would have had a lot more success in this class if I would have had access to the bot for labs. (Labs on average took me about 2 and a half hours to complete, and sometimes longer if I didn’t have access to a tutor).”
What students liked the least:
“Sometimes the answers were slightly vague. Of course, the bot can’t simply spit out the answer, but sometimes it was frustrating how it would say ‘you’re on the right track, but there seems to be a conceptual misunderstanding with __’ — the explanations for the blank could be a bit jargon-filled and didn’t always directly help me resolve the misunderstanding due to imprecise language.”
@@
Students generally appreciated 61A-Bot’s accessibility, debugging capabilities, and time savings. However, the hints were sometimes too vague for the students to make changes, while other times, the bot was too specific and gave away too much. Despite our attempts to steer GPT-4 towards rephrasing and a broader diversity of responses to repeat inputs, these incidents still occur.
4.3. Effects on Assignment Completion Time
@@
Having established that students like 61A-Bot, and that it can provide hints that appear useful, we began seeking ways to quantitatively identify what impact—if any—the bot was having on student outcomes. This proved to be a challenging task, in part because the impact is nuanced and specific, as we detail in this section. As mentioned earlier, our IRB does not allow differential access to tools in courses, so we could not run a randomized control trial. Instead, here we compare student performance from our semester of full deployment, Spring 2024 (SP24), with performance from prior semesters, going back through Spring 2021 (SP21), comprising a total of 1,643,613 data points from 6,034 students.
Though the Bot is deployed for homework assignments it is not available for lab assignments (which are similar in scope, but intended to be completed in groups) nor for projects (which are multi-week and build on prior results). To the extent that performance differences on homework assignments between pre- and post-deployment semesters are inconsistent with performance differences on lab assignments and projects, some of this difference may be attributable to the use of the Bot. In our course, content is typically sequenced as follows: new material is first introduced in lectures, then practiced in labs, reinforced in individual homework problems, and then applied to practice in projects.
4.3.1. Data
@@
Our course “autograder” records every attempt a student makes to test their code, storing a “snapshot” of student code on our instructional servers, along with a student identifier and a timestamp; we use these to reconstruct a coarse-grained history of student progress. Autograder use is unlimited, and students typically revise their programs repeatedly until they are ready to submit their final code. Because nearly all students submit code that successfully passes all test cases, the metric we consider here is time to completion, rather than passing test cases or other measures of code quality.
@@
We calculate an approximate total completion time by summing the timestamp deltas between snapshots, recognizing that we may miss the time spent on the problem before the first snapshot. To account for students completing the homework across multiple sessions, we ignore time deltas above a 60-minute threshold—a value chosen somewhat arbitrarily, but we confirm that the results we report here are robust to values in the range of 20 to 120 minutes.777We conducted an independent analysis looking only at snapshot counts rather than summing time deltas. Those results were consistent with the findings we report here, as expected from (Piech et al., 2012); we report times here because we find the time metric to be more straightforward to reason about.
We compare Spring 2024 completion times only to prior Spring semesters, because Fall-semester students at our institution typically have greater preparation than Spring-semester students (due in part to how CS major requirements are structured). Because problem repetition across homeworks and labs has some variation across semesters, we select a subset of problems to maximizes consistency, and report these aggregations as, e.g., “HW 1”. Occasionally, one assignment in a particular semester differs sufficiently from other semesters that we omit it entirely.
4.3.2. Analytical Method
We examine the distribution of student assignment completion times using Cumulative Distribution Function (CDF) plots for specific homework assignments, labs, and projects. This CDF can be read as “What fraction of students (y-axis) complete the homework in less than some number of minutes (x-axis)?” Because not all semesters used exactly the same problems in each assignment, we analyze only that subset of problems that appear identically across most semesters in a given assignment, summing those problems’ completion times and reporting the sum as the total homework completion time for the assignment. (Some semesters are missing from some plots because of this variability in problem inclusion for assignments as mentioned earlier.) As with the choice of completion time vs. submission counts, we opted for this method of aggregation and reporting because we found it to be the clearest presentation of the effects we observed.888We performed the same analysis on individual problems on their own, as well as by comparing individual semesters with their prior (Spring) semesters and comparing across different subsets of problems and prior years; the results we report here are robust to all of the variations we tried, including many others that we do not report on here for space.
@@ 4.3.3. Results
The data indicate that the Bot-available homework completion times in Spring 2024 are between faster among students in the middle percentile of completion times. This represents a reduction in time of Standard Deviations (SD) compared with completion times from prior Spring semesters, a large effect.
@@ 5. Discussion
Overall, our findings are consistent with a causal link between the availability of our bot and faster homework completion times. There is no substantially different teaching methodology, set of course instructors, or set of course materials in Spring 2024 that could account for the specific reduction in homework completion times without a similar reduction in lab completion times—to our knowledge (as course instructors), 61A-Bot itself is the major difference in Spring 2024, and the major difference between homeworks on the one hand, and labs and projects on the other.
The speedup differences among homeworks, projects, and labs, could be due to a few causal factors. First, some of the benefits of using the Bot on homeworks could transfer to projects, but not labs—perhaps because projects around a given topic are assigned after students have already had the Bot’s help on the corresponding homework, but not on the corresponding labs, which are assigned before the homework.
Second, students in Spring 2024 may simply have stronger prior preparation in a way that affects projects, but not labs—though we are not sure what kind of preparation that could be. The consistency of performance on the first few labs in Spring 2024 and prior semesters suggests that Spring 2024 students do not differ substantially in their prior preparation for CS1.
The apparent gradual increase in lab completion times (starting in lab 4, continuing in lab 8, see 3(b)) are worth noting because they do not correlate with a corresponding increase in homework completion time, suggesting that the availability of the bot may be a cause. Students could be relying on the bot for certain activities that they might otherwise have learned more effectively to perform on their own, such as debugging code on previously unseen topics. We find some possible corroborating evidence for this effect in the survey results, where a top request for the course is to make the bot available for lab assignments too (see §4.2). But the improvement in project completion times suggests that if there is over-reliance on the Bot in the lab context, it does not extend to the project context—and it could be that lab 8 is merely an outlier in this regard, and not indicative of a robust pattern.
5.1. Limitations
We have been careful to identify limitations throughout this report but include this section here for completeness. Most importantly, our study is entirely observational: we had limited levers for randomization, and there could be uncontrolled causes for the effects we observe—for example, it could be that students are using ChatGPT for their homework anyway, despite the availability of 61A-Bot. We have tried to account for as many of these factors as possible through mechanism-oriented reasoning (e.g., why would they use ChatGPT only for homeworks, but not labs?), but we can not prove causality with our study design.
@@ 6. Conclusions
Our results suggest that 61A-Bot helps students complete homework more quickly, with oversized impacts for students who spent the most time on homework—a benefit that might even disproportionately support goals towards inclusion in CS. But, ideally, this type of scaffolding should recede over time as learners become more confident (Soloway et al., 1994). 61A-Bot has not yet achieved this goal.
However, the news is not entirely bad: whatever skills students may come to depend on the bot scaffolding for—whether debugging, problem-solving, or something else—are skills we can try to teach explicitly in our course, now that students spend less time on homework. Of course, it may turn out that the best way for students to learn those skills is to struggle more in debugging their homework assignments on their own, but we are skeptical of this view. Fortunately, effective homework assistants, like our 61A-Bot, can also help future researchers start to disentangle these effects, leading to a better understanding of the pedagogical value of struggling to debug code.
Guidelines around the inclusion of AI-based course materials and tools suggest that these should only be incorporated when we have a good understanding that their benefits outweigh their costs (Bala et al., 2023). We believe that 61A-Bot clears this bar, but that further research into improving outcomes and mitigating the costs we have started to expose is critical and urgent.
@@ Acknowledgements.
- Denny et al. (2023b) Paul Denny, Juho Leinonen, James Prather, Andrew Luxton-Reilly, Thezyrie Amarouche, Brett A Becker, and Brent N Reeves. 2023b. Promptly: Using Prompt Problems to Teach Learners How to Effectively Utilize AI Code Generators. arXiv preprint arXiv:2307.16364 (2023).
- Denny et al. (2023c) Paul Denny, James Prather, Brett A Becker, James Finnie-Ansley, Arto Hellas, Juho Leinonen, Andrew Luxton-Reilly, Brent N Reeves, Eddie Antonio Santos, and Sami Sarsa. 2023c. Computing Education in the Era of Generative AI. arXiv preprint arXiv:2306.02608 (2023).
- Donlevy (2023) Katherine Donlevy. 2023. Harvard to roll out AI professors in flagship coding class for fall semester. https://nypost.com/2023/06/30/harvard-to-roll-out-ai-professors-in-flagship-coding-class-for-fall-semester/
@@
- Li et al. (2022) Yujia Li, David Choi, Junyoung Chung, Nate Kushman, Julian Schrittwieser, Rémi Leblond, Tom Eccles, James Keeling, Felix Gimeno, Agustin Dal Lago, et al. 2022. Competition-level code generation with alphacode. Science 378, 6624 (2022), 1092–1097.
@@
- Liu et al. (2024) Rongxin Liu, Carter Zenke, Charlie Liu, Andrew Holmes, Patrick Thornton, and David J. Malan. 2024. Teaching CS50 with AI: Leveraging Generative Artificial Intelligence in Computer Science Education. In Proceedings of the 55th ACM Technical Symposium on Computer Science Education V. 2 (Portland, OR, USA) (SIGCSE 2024). Association for Computing Machinery, New York, NY, USA, 1927.
@@
- Microsoft (2024) Microsoft. 2024. Azure OpenAI. https://azure.microsoft.com/en-us/products/ai-services/openai-service/
@@
- Zamfirescu-Pereira et al. (2023) J.D. Zamfirescu-Pereira, Heather Wei, Amy Xiao, Kitty Gu, Grace Jung, Matthew G Lee, Bjoern Hartmann, and Qian Yang. 2023. Herding AI Cats: Lessons from Designing a Chatbot by Prompting GPT-3. In Proceedings of the 2023 ACM Designing Interactive Systems Conference (Pittsburgh, PA, USA).

### v2

@@ 4.2. Student Reception
Student feedback suggests that students also found the Bot helpful. We solicited this feedback in two ways: First, we queried students for their assessment of each individual hint, which we received for approximately 27% () of queries.777These counts reflect queries from only those students who gave consent for their data to be used for research, and only during the CS1 portion of our course. Of these, 70% () were rated as “helpful,” with 45% of those, , reporting that the problem was now resolved. A further 10% () were rated as “not helpful, but made sense,” while the final 20% () were rated as insensible, misleading, or wrong.
@@
In Spring 2024, we conducted a non-anonymous survey at the end of the semester, to which 89% () of students responded. On a scale from 1 to 5, we asked students to rate their bot usage, bot helpfulness, bot reliability, and overall satisfaction with the bot. Finally, we asked them whether or not they recommend that the bot be available to students in future semesters. The results from these surveys can be found in Figure 2.
Note that these results include both our partial- (Fall 2023) and full-deployment (Spring 2024) semesters; the Fall 2023 results thus reflect usage only post-deployment. Additionally, these surveys were non-anonymous to track individual participation (for pedagogical goals independent of this project); because of this non-anonymity, however, we did not ask whether students relied on ChatGPT or other prohibited (by course policy) methods of support, reasoning that we were unlikely to be able to rely on such results.
@@
In the Spring 2024 survey, we also asked two optional free-response questions: What did students like the most and least about learning with the bot? We include a representative response for each here:
@@
Most liked: “What I loved about the bot is that it allowed me to get feedback when I didn’t have access to a tutor. Accordingly, instead of banging my head against the wall for hours, I was able to get feedback about what I was doing wrong and correct the mistakes. For me personally, I would have had a lot more success in this class if I would have had access to the bot for labs. (Labs on average took me about 2 and a half hours to complete, and sometimes longer if I didn’t have access to a tutor).” Least liked: “Sometimes the answers were slightly vague. Of course, the bot can’t simply spit out the answer, but sometimes it was frustrating how it would say ‘you’re on the right track, but there seems to be a conceptual misunderstanding with __’ — the explanations for the blank could be a bit jargon-filled and didn’t always directly help me resolve the misunderstanding due to imprecise language.”
@@
Students generally appreciated 61A Bot’s accessibility, debugging capabilities, and time savings. However, the hints were sometimes too vague for the students to make changes, while other times, the bot was too specific and gave away too much. Despite our attempts to steer GPT-4 towards rephrasing and a broader diversity of responses to repeat inputs, these incidents still occur.
4.3. Effects on Homework
@@
To understand the effects of 61A Bot on students in our course, we compare student performance from our semester of full deployment, Spring 2024 (SP24), with performance from prior Spring semesters (see §3.1), going back through Spring 2021 (SP21)—comprising a total of 1,643,613 data points from 6,034 students. We chose this method of analysis in large part because our IRB does not allow differential access to tools in courses, and thus we could not run a randomized control trial.
Though 61A Bot is available to students for homework assignments, it is not available for lab assignments nor for projects (see §3.1). To the extent that performance differences on homework assignments between pre- and post-deployment semesters are inconsistent with performance differences on lab assignments and projects, some of this difference may be attributable to the use of the Bot. Though course content is substantially similar across the semesters we report on here, there is some variation in course staff, individual lectures, and specific problems within labs and homework assignments—where homework and lab problems have changed more than trivially, we have omitted them from our comparisons, examining only those that are identical in nature and sequencing within our course. (We confirm, too, that the specific assignments we report on here are representative of the full set of assignments, and we are aware of no other major changes in course content, delivery, or staffing occurred during this time.)
4.3.1. Data & Analytical Method
@@
Our course “autograder” records every attempt a student makes to test their code, storing a “snapshot” of student code on our instructional servers, along with a student identifier and a timestamp; we use these to reconstruct a history of student progress. Autograder use is unlimited, and students typically revise their programs repeatedly until they are ready to submit their final code. Nearly all students submit code that successfully passes all test cases, so examining the final submitted code artifact alone does not necessarily provide useful insight into impacts. Thus the metric we consider here is time to completion, rather than passing test cases or other measures of code quality.
@@
We calculate an approximate total completion time by summing the timestamp deltas between snapshots. To account for students completing the homework across multiple sessions, we ignore time deltas above a 60-minute threshold—a value chosen somewhat arbitrarily, but we confirm that the results we report here are robust to values in the range of 20 to 120 minutes.888We also conducted an independent analysis looking only at snapshot counts rather than summing time deltas. Those results were consistent with the findings we report here, as expected from (Piech et al., 2012); we report times here because we find the time metric to be more straightforward to reason about. For clarity, we aggregate individual repeated problems into “assignments,” e.g., “HW 1.” Occasionally, one assignment in a particular semester differs sufficiently from other semesters that we omit it entirely.999Like our choice of completion time vs. submission counts, we opted for this method of aggregation because we found it to be the clearest presentation of the effects we observed. We performed the analysis described in this section with individual problems on their own, as well as by comparing individual semesters with their prior (Spring) semesters and comparing across different subsets of problems and prior years; the results we report here are robust to all of the variations we tried, including many others that we do not report on here for space.
We then report the distribution of student assignment completion times using Cumulative Distribution Function (CDF) plots for specific homework assignments, labs, and projects. This CDF can be read as “What fraction of students (y-axis) complete the homework in less than some number of minutes (x-axis)?”, capturing how long students take to complete these assignments across the full distribution of completion times.
@@ 4.3.2. Results
The data indicate that the (Bot-available) homework completion times in Spring 2024 are between faster among students in the middle percentile of completion times. This represents a reduction in time of Standard Deviations (SD) compared with completion times from prior Spring semesters, a large effect.
@@ 5. Discussion
Overall, our findings are consistent with a causal link between the availability of our bot and faster homework completion times. In the context of our course (see §4.3.1), reduced completion times are likely to be a result of students reaching a correct solution more quickly, rather than stopping before reaching a correct solution, or making other mistakes, as might be the case in domains where assignments have a greater variety of possible outcomes.
Though we did not carefully examine (and don’t make claims about) student learning outcomes here, we have reason to believe that students are not performing dramatically worse after deployment. In particular, though student exam performance is sufficiently inconsistent (due to dramatic exam coverage differences from term to term) that we did not report on it here, it does not appear that Spring 2024 students performed substantially worse than in prior terms. Additionally, the differences in performance on assignments where the Bot was not available (labs and projects) were much smaller. We can’t say for certain, but we have not found much evidence for a major decrease in outcomes.
If indeed this reduction in homework time results in little or no learning loss, we can then ask: What implications does this have for CS1 courses? Are there other costs unrelated to learning loss? And what other benefits might accrue?
On the costs side, with most homework help requests going through an LLM, our human TAs may struggle more to stay on top of common challenges among students, leading to a looser feedback loop. However, this challenge could be addressed with new tools that help TAs aggregate over automated help requests—but now with LLM-generated hints and student feedback on whether those hints are helpful. That is, of course, assuming CS1 courses do not elect to reduce staff in response.
Given the mode of access of 61A Bot, one likely change in student learning is a reduction in the ability to read and comprehend error traces. Now that our students have an easy way to get natural language feedback on failed test cases, they are much less incentivized to try to understand why a test case failed by reading the trace.
On the benefits side, instructors could use this extra student time to cover more material, such as effective debugging techniques or reading stack traces. Or, students could continue to simply spend less time on the course. Similarly, TAs could spend less time debugging straightforward homework errors and more time focused on other forms of support.
These factors all point to a need for more research to better understand the actual costs and to inform decisions made in the hope of realizing actual benefits beyond saved student and TA time.
5.1. Limitations & Threats to Validity
Differences in prior preparation could explain why students complete homework more quickly in Spring 2024—but any such differences would also have to explain the lack of a decrease in lab completion times. In fact, the consistency in early lab completion times over the 4 semesters we examined suggests that our student populations do not differ significantly in prior ability.
Similarly, differences in course content delivery, staffing, structure, etc., would be expected to impact both labs and homeworks, as lectures and discussion sections for given topics come before labs and homeworks. To our knowledge, no additional homework support was provided in Spring 2024 beyond 61A Bot—no additional hints or support unique to Spring 2024 were offered in lecture or group sessions.
Finally, our study is entirely observational: we had limited levers for randomization, and there could be uncontrolled causes for the effects we observe—for example, it could be that students are using ChatGPT for their homework or projects anyway, despite the prohibition on ChatGPT use and the availability of 61A Bot. ChatGPT was originally released in November 2022, but we observed no similar effect in Spring 2023 compared with prior terms.
@@ 6. Conclusions
Our results suggest that 61A Bot reduces demands on staff and helps students complete homework more quickly, with oversized impacts for students who spent the most time on homework—a benefit that might even disproportionately support goals towards inclusion in CS. But, ideally, this type of scaffolding should recede over time as learners become more confident (Soloway et al., 1994). 61A Bot has not yet clearly achieved this goal.
Guidelines around the inclusion of AI-based course materials and tools suggest that these should only be incorporated when we have a good understanding that their benefits outweigh their costs (Bala et al., 2023). Yet even if the primary outcome of 61A Bot is limited to a reduction in homework completion time with no other benefits, we believe that 61A Bot clears this bar—but that further research into improving outcomes and mitigating the costs we have started to expose is critical and urgent.
@@ Acknowledgements.
- Bassner et al. (2024) Patrick Bassner, Eduard Frankford, and Stephan Krusche. 2024. Iris: An AI-Driven Virtual Tutor for Computer Science Education. In Proceedings of the 2024 on Innovation and Technology in Computer Science Education V. 1 (Milan, Italy) (ITiCSE 2024). Association for Computing Machinery, New York, NY, USA, 394–400. https://doi.org/10.1145/3649217.3653543
@@
- Denny et al. (2024) Paul Denny, Stephen MacNeil, Jaromir Savelka, Leo Porter, and Andrew Luxton-Reilly. 2024. Desirable Characteristics for AI Teaching Assistants in Programming Education. In Proceedings of the 2024 on Innovation and Technology in Computer Science Education V. 1 (Milan, Italy) (ITiCSE 2024). Association for Computing Machinery, New York, NY, USA, 408–414. https://doi.org/10.1145/3649217.3653574
- Denny et al. (2023b) Paul Denny, James Prather, Brett A Becker, James Finnie-Ansley, Arto Hellas, Juho Leinonen, Andrew Luxton-Reilly, Brent N Reeves, Eddie Antonio Santos, and Sami Sarsa. 2023b. Computing Education in the Era of Generative AI. arXiv preprint arXiv:2306.02608 (2023).
@@
- Liu et al. (2024a) Rongxin Liu, Carter Zenke, Charlie Liu, Andrew Holmes, Patrick Thornton, and David J. Malan. 2024a. Teaching CS50 with AI: Leveraging Generative Artificial Intelligence in Computer Science Education. In Proceedings of the 55th ACM Technical Symposium on Computer Science Education V. 1 (Portland, OR, USA) (SIGCSE 2024). Association for Computing Machinery, New York, NY, USA, 750–756. https://doi.org/10.1145/3626252.3630938
- Liu et al. (2024b) Rongxin Liu, Carter Zenke, Charlie Liu, Andrew Holmes, Patrick Thornton, and David J. Malan. 2024b. Teaching CS50 with AI: Leveraging Generative Artificial Intelligence in Computer Science Education. In Proceedings of the 55th ACM Technical Symposium on Computer Science Education V. 2 (Portland, OR, USA) (SIGCSE 2024). Association for Computing Machinery, New York, NY, USA, 1927.
@@
- Prather et al. (2024) James Prather, Juho Leinonen, Natalie Kiesler, Jamie Gorson Benario, Sam Lau, Stephen MacNeil, Narges Norouzi, Simone Opel, Virginia Pettit, Leo Porter, et al. 2024. How Instructors Incorporate Generative AI into Teaching Computing. In Proceedings of the 2024 on Innovation and Technology in Computer Science Education V. 2. 771–772.